Hayhayz (hayhayz):

Larry and Peggy are making decisions about their bank accounts. Larry wants to deposit $350 as a principle amount, with an interest of 4% compounded quarterly. Peggy wants to deposit $350 as the principle amount, with an interest of 6% compounded monthly. Explain which method results in more money after 2 years. Show all work.

OpenStudy (anonymous):

A = P(1+r/n)^nt is the formula for compound interest. Plug in the values. n is 12

OpenStudy (anonymous):

So A = P (1 + 4/12)^12(35)?

OpenStudy (anonymous):

350*

OpenStudy (anonymous):

hey i said it first

OpenStudy (mathstudent55):

You don't need to calculate anything to answer this question. If all else is the same, a higher interest rate yields more interest. If all else is equal, compounding interest more often yields more interest. In this case, one choice has both a higher interest rate and compounding interest more often. It must yield more interest at the end of the same amount of time.

OpenStudy (mathstudent55):

\(F = P(1 + \dfrac{r}{n})^{nt} \) Larry: \(F = P(1 + \dfrac{r}{n})^{nt} = $350(1 + \dfrac{0.04}{4})^{(4)(2)}\) Peggy: \(F = P(1 + \dfrac{r}{n})^{nt} = $350(1 + \dfrac{0.06}{12})^{(12)(2)}\)